Giant Film Company just paid a dividend of $3 per share yesterday and has an EPS of $3 and a required return of 10 percent per year. Please answer the following questions: a. If Giant Film’s future dividends are expected to stay at $3 forever, what’s Giant Film’s current stock price? b. If Giant Film’s future dividends are expected to grow at 5% per year, what’s Giant Film’s current stock price?
Answer #1

a. price = dividend/required rate = 3/0.1 = 30

b.

As per DDM
Price = recent dividend* (1 + growth rate )/(cost of equity - growth rate)
Price = 3 * (1+0.05) / (0.1 - 0.05)
Price = 63